OpenStudy (anonymous):

What is the inverse of this quadratic function? y=4-x^2+3x,,, thx

OpenStudy (anonymous):

just plug y in for x and x in for y; then solve for y

OpenStudy (anonymous):

x=4-y^2+3y x=-y^2+3y+4 0=-y^2+3y-x+4 But now what?

OpenStudy (anonymous):

solve for y now... from that last line, can you solve for y? that will be your inverse. but also consider the domain for f^-1(x) now....